OpenStudy (anonymous):

Combine like terms, show work?! Help?! 2x – 5 + 6x + 3 = 14

OpenStudy (anonymous):

Just add (2x+6x)+(-5+3)=14 Then after you're done, leave x on the left side, and put numbers on the right side, (2x+6x) = 14 - (-5+3)